OpenStudy (anonymous):

George pushes a wheelbarrow for a distance of 12 meters at a constant speed for 35 seconds by applying a force of 20 newtons. What is the power applied to push this wheelbarrow?

OpenStudy (anonymous):

\[\color{green}{P = \frac{W}{t}\\ = \frac{Fd}{t}\\ = \frac{20~\text{N} \times 12~\text{m}}{35~\text{s} \\ = \underline{~~~~~~~~~}~ \text{W}}}\]